OpenStudy (anonymous):

Given the function g(x) = −4x + 5, find g(−4).

OpenStudy (cwrw238):

replace the x in -4x + 5 with -4 and simplify

OpenStudy (anonymous):

just put x=-4 i the function g(-4)=-4(-4)+5 g(-4)=16+5 g(-4)=21